# Fundamental role of Dirac supersingletons in particle theory

### v1

@@ Abstract
As shown in the famous Dyson’s paper ”Missed Opportunities”, even from purely mathematical considerations (without any physics) it is clear that Poincare quantum symmetry is a special degenerate case of de Sitter quantum symmetries. Then the question arises why in particle physics Poincare symmetry works with very high accuracy. The usual answer to this question is that a theory in de Sitter space becomes a theory in Minkowski space in the formal limit when the radius of de Sitter space tends to infinity. However, de Sitter and Minkowski spaces are purely classical concepts, and in quantum theory the answer to this question must be given only in terms of quantum concepts. At the quantum level, Poincare symmetry is a good approximate symmetry if the eigenvalues of the representation operators $M_{4\mu}$ of the anti-de Sitter algebra are much greater than the eigenvalues of the operators $M_{\mu\nu}$ ($\mu,\nu=0,1,2,3$). We show that an explicit solution with such properties exists within the framework of the approach proposed by Flato and Fronsdal where standard elementary particles are bound states of two Dirac singletons.
@@ 1 Problem statement
As noted above, symmetry at the quantum level should be defined in terms of the the corresponding Lie algebras, and in [2], the statements a)-c) have been reformulated in such terms. It has also been shown that the fact that quantum theory is more general than classical theory follows even from purely mathematical considerations because formally the classical symmetry algebra can be obtained from the symmetry algebra in quantum theory by contraction $\hbar\to 0$. For these reasons, the most general description in terms of ten-dimensional Lie algebras should be carried out in terms of quantum dS or AdS symmetry.
@@
At the classical (non-quantum) level, the transition from dS or AdS symmetry to Poincare one is explained as follows. When the radius $R$ of dS or AdS space becomes infinitely large, the angular momentum $M$ of a particle moving in this space also becomes infinitely large. In the formal limit $R\to\infty$ when dS or AdS space transform into flat Minkowski space, the motion of a particle in such space must be described by momentum $p=M/R$ which is finite in this limit.
One can raise the question why Poincare symmetry works with high accuracy in particle physics. At the classical level, the explanation of this fact is that we live in dS or AdS space whose radius $R$ is very large. However, as noted above, the concept of background space is purely classical. Therefore, the question arises whether the answer to the above question can be given within the framework of purely quantum theory, without involving classical concepts. As follows from the above definition of contraction from dS or AdS algebra to Poincare one, Poincare invariance works with high accuracy, provided that in nature such states are realized in which the eigenvalues of the operators $M^{4\mu}$ are much greater than the eigenvalues of the operators $M^{\mu\nu}$ ($\mu,\nu$=0,1,2,3). In this paper, we propose a scenario that describes such a situation.
The paper is organized as follows. In Sec. 2 we explain why supersymmetric AdS symmetry is more general (fundamental) than standard AdS symmetry. In Sec. 3 we explicitly describe the construction of the IR for Dirac supersingletons and explicit relations between representation operators in this IR and representation operators of the AdS algebra. Then in Sec. 4 it is explicitly shown that there exists a scenario when Poincare symmetry works with a high accuracy.
@@ 2 Supersymmetry
Note that representations of the standard Poincare superalgebra are described by 14 operators. Ten of them are the representation operators of the Poincare algebra—four momentum operators and six operators of the Lorentz algebra, and in addition, there are four fermionic operators. The anticommutators of the fermionic operators are linear combinations of the Lorentz algebra operators, the commutators of the fermionic operators with the Lorentz algebra operators are linear combinations of the fermionic operators and the fermionic operators commute with the momentum operators. However, the latter are not bilinear combinations of fermionic operators.
From the formal point of view, representations of the AdS so(2,3) superalgebra osp(1,4) are also described by 14 operators — ten representation operators of the so(2,3) algebra and four fermionic operators. There are three types of relations: the operators of the so(2,3) algebra commute with each other as usual (see Eqs. (2)), anticommutators of the fermionic operators are linear combinations of the so(2,3) operators and commutators of the latter with the fermionic operators are their linear combinations. However, in fact, representations of the osp(1,4) superalgebra can be described exclusively in terms of the fermionic operators. The matter is as that in the general case, the anticommutators of four operators form ten independent linear combinations. Therefore, ten bosonic operators can be expressed in terms of fermionic ones. This is not the case for the Poincare superalgebra since the Poincare algebra operators are obtained from the so(2,3) ones by contraction. One can say that the representations of the osp(1,4) superalgebra is an implementation of the idea that supersymmetry is the extraction of the square root from the usual symmetry (by analogy with the treatment of the Dirac equation as a square root from the Klein-Gordon equation).
@@ 3 Dirac supersingleton
If $m$ is the mass of the particle in Poincare invariant theory then its mass $\mu$ in AdS theory is dimensionless and the relation between $\mu$ and $m$ is $\mu=mR$ where $R$ is the contraction parameter for the transition from AdS to Poincare symmetry. As explained in [4], $R$ has nothing to do with the radius of classical AdS space, $R$ is fundamental to the same extent as $c$ and $\hbar$, and the problem why the value of $R$ is as is does not arise. The data on cosmological acceleration show that, at the present stage of the universe, $R$ is of the order of $10^{26}m$ [4]. Therefore, even for elementary particles, the AdS masses are very large. For example, the AdS masses of the electron, the Earth and the Sun are of the order of $10^{39}$, $10^{93}$ and $10^{99}$, respectively. The fact that even the AdS mass of the electron is so large might be an indication that the electron is not a true elementary particle. In addition, the present upper level for the photon mass is $10^{-17}ev$ or less. This value seems to be an extremely tiny quantity. However, the corresponding AdS mass is of the order of $10^{16}$ and so, even the mass which is treated as extremely small in Poincare invariant theory might be very large in AdS invariant theory.
As shown in [4], in standard quantum theory, dS symmetry is more general than AdS one but in the framework of this symmetry it is not possible to describe neutral elementary particles, i.e., particles which are equivalent to the their antiparticles. In FQT, dS and AdS symmetries are equivalent and, as shown in [2], in this theory also there are no neutral elementary particles. In particular, even the photon is not elementary.
@@
There exists a wide literature (see e.g. [9, 10] and references therein) where this problem is investigated from the point of view of standard AdS QFT. For example, in AdS QFT, singleton fields live on the boundary at infinity of the AdS bulk (boundary which has one dimension less than the bulk). However, as noted in Sec. 1, the explanation in the framework of quantum theory should not involve classical spaces.
@@
As argued in [2], in FQT, the properties of Dirac singletons are even more remarkable than in standard theory. Here the eigenvalues of the operators $h_{1}$ and $h_{2}$ for singletons in FQT are $(p+1)/2,(p+3)/2,(p+5)/2...$, i.e. huge numbers if $p$ is huge. Hence Poincare limit for Dirac singletons in FQT has no physical meaning and they cannot be observable.
@@
4 Why Poincare symmetry in particle theory works with a high accuracy
@@
Even when we work in FQT and consider states of supersingletons in which the quantum numbers $j,k$ are much less than $p$, then, with high accuracy, we can apply standard mathematics. We assume that, although the numbers $j,k$ can be very large, they are still much less than $p$. Therefore, in what follows we consider Dirac supersingletons only within the framework of standard mathematics.
@@
We now treat $(d_{1}^{\prime},d_{2}^{\prime},d_{1}^{\prime\prime},d_{2}^{\prime\prime})$ as the operators in the Hilbert space related by Hermitian conjugation as $(d_{1}^{\prime})^{*}=d_{1}^{\prime\prime}$ and $(d_{2}^{\prime})^{*}=d_{2}^{\prime\prime}$. Then, as follows from Eq. (6), the squared norm of the element $e(j,k)$ is equal to
@@
Therefore, the normalized basis vectors can be defined as
@@
We consider the supersingleton wave functions $\sum_{jk}c(j,k){\tilde{e}}(j,k)$ in the framework of semiclassical approximation when the coefficients $c(j,k)$ are not equal to zero only at $j\in(j_{1},j_{2}),\,\,k\in(k_{1},k_{2})$ where $j_{2}-j_{1}\ll j_{1},\,\,k_{2}-k_{1}\ll k_{1}$ and the values of $|c(j,k)|$ at such $j,k$ are approximately the same. We define the angular dependence of the coefficients as $c(j,k)=|c(j,k)|exp[i(j+k)\chi+i(j-k)\varphi]$. Then taking into account Eqs. (4,6,8), and the definition of the basis elements and the coefficients $c(j,k)$, direct calculation shows that, in semiclassical approximation, the operators $M_{ab}$ can be replaced by their numerical values:
@@
$M_{10}=jsin(2\varphi+2\chi)+ksin(2\varphi-2\chi)$
@@
$M_{24}=jsin(2\varphi+2\chi)-ksin(2\varphi-2\chi),\,\,M_{04}=j+k+1$ (9)
@@
Let us now consider a system of two free supersingletons 1 and 2. The AdS superalgebra representation for such a system is the tensor product of the representations for supersingletons 1 and 2, and the representation operators are the sums of the corresponding operators: $M_{ab}=M_{ab}^{(1)}+M_{ab}^{(2)}$. The question arises under what conditions a system consisting of such supersingletons can be described in the framework of Poincare symmetry with a high accuracy. As noted above, a necessary condition for this is that for the system as a whole, the eigenvalues of the operators $M_{4\mu}$ must be much greater than the eigenvalues of the operators $M_{\mu\nu}$.

### v2

@@ Abstract
As shown in the famous Dyson’s paper ”Missed Opportunities”, even from purely mathematical considerations (without any physics) it follows that Poincare quantum symmetry is a special degenerate case of de Sitter quantum symmetries. Then the question arises why in particle physics Poincare symmetry works with a very high accuracy. The usual answer to this question is that a theory in de Sitter space becomes a theory in Minkowski space in the formal limit when the radius of de Sitter space tends to infinity. However, de Sitter and Minkowski spaces are purely classical concepts, and in quantum theory the answer to this question must be given only in terms of quantum concepts. At the quantum level, Poincare symmetry is a good approximate symmetry if the eigenvalues of the representation operators $M_{4\mu}$ of the anti-de Sitter algebra are much greater than the eigenvalues of the operators $M_{\mu\nu}$ ($\mu,\nu=0,1,2,3$). We show that explicit solutions with such properties exist within the framework of the approach proposed by Flato and Fronsdal where particles that are considered elementary in the standard theory are bound states of two Dirac singletons.
@@ 1 Problem statement
As noted above, symmetry at the quantum level should be defined in terms of Lie algebras, and in [2], the statements a)-c) have been reformulated in such terms. It has also been shown that the fact that quantum theory is more general than classical theory follows even from purely mathematical considerations because formally the classical symmetry algebra can be obtained from the symmetry algebra in quantum theory by contraction $\hbar\to 0$. For these reasons, the most general description in terms of ten-dimensional Lie algebras should be carried out in terms of quantum dS or AdS symmetry.
@@
At the classical (non-quantum) level, the transition from dS or AdS symmetry to Poincare one is explained as follows. When the radius $R$ of dS or AdS space becomes infinitely large, the angular momentum $M$ of a particle moving in this space also becomes infinitely large. In the formal limit $R\to\infty$ when dS or AdS space transforms into flat Minkowski space, the motion of a particle in such space must be described by momentum $p=M/R$ which is finite in this limit.
One can raise the question why Poincare symmetry works with high accuracy in particle physics. At the classical level, the explanation of this fact is that we live in dS or AdS space whose radius $R$ is very large. However, as noted above, the concept of background space is purely classical. Therefore, the question arises whether the answer to the above question can be given within the framework of purely quantum theory, without involving classical concepts. As follows from the above definition of contraction from dS or AdS algebra to Poincare one, Poincare invariance works with a high accuracy, provided that in nature such states play the major role in which the eigenvalues of the operators $M^{4\mu}$ are much greater than the eigenvalues of the operators $M^{\mu\nu}$ ($\mu,\nu$=0,1,2,3). In this paper, we propose a scenario that describes such a situation.
The paper is organized as follows. In Sec. 2 we explain why supersymmetric AdS symmetry is more general (fundamental) than standard AdS symmetry. In Sec. 3 we describe how the CPT transformation works at the quantum level. In Sec. 4 we explicitly describe the construction of the IR for Dirac supersingletons and explicit relations between representation operators in this IR and representation operators of the AdS algebra. Then in Sec. 5 it is explicitly shown that there exist scenarios when Poincare symmetry works with a high accuracy.
@@ 2 Supersymmetry
From the formal point of view, representations of the AdS so(2,3) superalgebra osp(1,4) are similar to representations of the Poincare superalgebra because they are also described by 14 operators — ten representation operators of the so(2,3) algebra and four fermionic operators. There are three types of relations: the operators of the so(2,3) algebra commute with each other as usual (see Eqs. (2)), anticommutators of the fermionic operators are linear combinations of the so(2,3) operators and commutators of the latter with the fermionic operators are their linear combinations. However, in fact, representations of the osp(1,4) superalgebra can be described exclusively in terms of the fermionic operators. The matter is as that in the general case, the anticommutators of four operators form ten independent linear combinations. Therefore, ten bosonic operators can be expressed in terms of fermionic ones. This is not the case for the Poincare superalgebra since the Poincare algebra operators are obtained from the so(2,3) ones by contraction. One can say that the representations of the osp(1,4) superalgebra is an implementation of the idea that supersymmetry is the extraction of the square root from the usual symmetry (by analogy with the treatment of the Dirac equation as a square root from the Klein-Gordon equation).
@@
3 CPT transformation in osp(1,4) invariant theory
In Poincare invariant particle theory, the CPT transformation is considered the most general discrete spacetime transformation. Based on what was said in Sec. 1, at the quantum level, this transformation should be considered not from the point of view of Minkowski space, but at the operator level. We use $\theta$ to denote the operator corresponding to the classical CPT transformation. As Wigner noted [6], since the sign of the energy must remain positive under the $\theta$ transformation, the operator $\theta$ must be not unitary, but antiunitary, that is, it can be represented as $\theta=\beta K$ where $\beta$ is a unitary operator, and $K$ is the complex conjugation operator. As shown by Schwinger [7], the problem of the sign of energy can also be solved if instead of the antiunitary transformation the transpose operation is used. In this work we use Wigner’s approach.
As shown by Wigner [6] (see also [8]), transformation $\theta$ transforms the representation operators of the Poincare algebra in Eq. (1) as
$\theta P_{\mu}\theta^{-1}=P_{\mu},\quad\theta M_{\mu\nu}\theta^{-1}=-M_{\mu\nu}$ (6)
The question arises of how to generalize these relationships to the case of dS and AdS theories and, as noted in Sec. 1, this generalization should not involve dS and AdS spaces. The issue of CPT transformation in such theories has been considered by many authors. However, to the best of our knowledge, these authors considered the CPT transformation only from the point of view of transformations of fields on the dS and AdS spaces and did not consider a direct generalization of Eq. (6). Moreover, as noted in Sec. 2, the superalgebra osp(1,4) is a generalization of the algebra so(2,3) to the case of supersymmetry, and, to the best of our knowledge, Eqs. (6) have not been generalized to representations of this superalgebra.
For this purpose, it is necessary to define how the operator $\theta$ transforms the operators $(d_{1}^{\prime},d_{2}^{\prime},d_{1}^{\prime\prime},d_{2}^{\prime\prime})$. We define such a transformation as follows:
$\theta d_{1}^{\prime}\theta^{-1}=-id_{2}^{\prime},\,\,\theta d_{2}^{\prime}\theta^{-1}=id_{2}^{\prime},\,\,\theta d_{1}^{\prime\prime}\theta^{-1}=id_{2}^{\prime\prime},\,\,\theta d_{2}^{\prime\prime}\theta^{-1}=-id_{1}^{\prime\prime}$ (7)
It is easy to see that the second of these relations follows from the first, and the fourth follows from the third, taking into account the fact that the operator $\theta$ is antiunitary.
Now, based on Eqs. (4) and (7) we conclude that
$\theta M_{4\mu}\theta^{-1}=M_{4\mu},\quad\theta M_{\mu\nu}\theta^{-1}=-M_{\mu\nu},\quad\mu,\nu=0,1,2,3$ (8)
and this is a generalization of Eq. (6) to the case of representations of the algebra so(2,3) because, as noted in Sec. 1, when contracting representations of the algebra AdS into representations of the Poincare algebra, the operators $M_{\mu\nu}$ are not affected, and the operators $M_{4\mu}$ go into $P_{\mu}$. This result is also natural from the observation that, as it is easy to see, Eqs. (2) are invariant under substitutions
$M_{4\mu}\to M_{4\mu},\quad M_{\mu\nu}\to-M_{\mu\nu},\quad i\to-i$
That these substitutions involve $i\to-i$ follows from the fact that the operator $\theta$ is antiunitary.
@@ 4 Dirac supersingleton
If $m$ is the mass of a particle in Poincare invariant theory then its mass $\mu$ in AdS theory is dimensionless and the relation between $\mu$ and $m$ is $\mu=mR$ where $R$ is the contraction parameter for the transition from AdS to Poincare symmetry. As explained in [4], $R$ has nothing to do with the radius of classical AdS space, $R$ is fundamental to the same extent as $c$ and $\hbar$, and the problem why the value of $R$ is as is does not arise. The data on cosmological acceleration show that, at the present stage of the universe, $R$ is of the order of $10^{26}m$ [4]. Therefore, even for elementary particles, the AdS masses are very large. For example, the AdS masses of the electron, the Earth and the Sun are of the order of $10^{39}$, $10^{93}$ and $10^{99}$, respectively. The fact that even the AdS mass of the electron is so large might be an indication that the electron is not a true elementary particle. In addition, the present upper level for the photon mass is $10^{-17}ev$ or less. This value seems to be an extremely tiny quantity. However, the corresponding AdS mass is of the order of $10^{16}$ and so, even the mass which is treated as extremely small in Poincare invariant theory might be very large in AdS invariant theory.
As shown in [4], in standard quantum theory, dS symmetry is more general than AdS one but in the framework of dS symmetry it is not possible to describe neutral elementary particles, i.e., particles which are equivalent to the their antiparticles. In FQT, dS and AdS symmetries are equivalent and, as shown in [2], in this theory also there are no neutral elementary particles. In particular, even the photon is not elementary.
@@
There exists a wide literature (see e.g., [12, 13] and references therein) where this problem is investigated from the point of view of standard AdS QFT. For example, in AdS QFT, singleton fields live on the boundary at infinity of the AdS bulk (boundary which has one dimension less than the bulk). However, as noted in Sec. 1, the explanation in the framework of quantum theory should not involve classical spaces.
@@
As argued in [2], in FQT, the properties of Dirac singletons are even more remarkable than in standard theory. Here the eigenvalues of the operators $h_{1}$ and $h_{2}$ for singletons in FQT are $(p+1)/2,(p+3)/2,(p+5)/2...$, i.e., huge numbers if $p$ is huge. Hence Poincare limit for Dirac singletons in FQT has no physical meaning and they cannot be observable.
@@
5 Why Poincare symmetry in particle theory works with high accuracy
@@
Let us first consider the case of neutral particles, which are considered elementary in the standard theory. As noted in Sec. 4, the standard theory has problems describing such particles, and in FQT neutral particles cannot be elementary. Therefore, based on the comments in Sec. 4, it is natural to consider the case when such particles are bound states of two sigletons. In this case, it is natural to treat the sigletons in the bound state are antiparticles for each other.
Therefore, if the first singleton is considered a particle for which representations of the AdS algebra are described by the operators (2), then the second singleton is naturally considered an antiparticle for which representations of the AdS algebra are described by the operators (2) transformed using the operator $\theta$, i.e., by the operators (8).
The osp(1,4) superalgebra representation for a system of two supersingletons is the tensor product of the representations for supersingletons 1 and 2, and the representation operators are the sums of the corresponding operators: $M_{ab}=M_{ab}^{(1)}+M_{ab}^{(2)}$. This means that if $\Psi_{1}$ is the state of the first supersingleton, and $\Psi_{2}$ is the state of the second supersingleton, which is the antiparticle for the first, then the operator $M_{ab}$ acts on the tensor product of these supersingletons as
$M_{ab}(\Psi_{1}\times\Psi_{2})=(M_{ab}^{(1)}\Psi_{1}\times\Psi_{2})+(\Psi_{1}\times M_{ab}^{(2)}\Psi_{2})$ (10)
The question arises under what conditions a system consisting of such supersingletons can be described in the framework of Poincare symmetry with a high accuracy. As noted above, a necessary condition for this is that for the system as a whole, the eigenvalues of the operators $M_{4\mu}$ must be much greater than the eigenvalues of the operators $M_{\mu\nu}$. Since the operators $M_{ab}^{(1)}$ and $M_{ab}^{(2)}$ are selfadjoined, all their eigenvalues are real. As follows from Eq. (8), if $M_{ab}^{(2)}=\theta M_{ab}^{(1)}\theta^{-1}$ then the operators $M_{4\mu}^{(1)}$ and $M_{4\mu}^{(2)}$ have the same set of eigenvalues and the sets of eigenvalues for the operators $M_{\mu\nu}^{(1)}$ and $M_{\mu\nu}^{(2)}$ differ by sign.
Also, the question arises what the states of two supersingletons must be to combine them into one particle, which is considered elementary in the standard theory. Since a system of two supersingletons is described by the tensor product of single supersingleton representations, then, from the point of view of standard theory, supersingletons do not interact with each other, but, as noted in Sec. 4, in FQT they actually interact. Nevertheless, it is natural to think that supersigletons in a bound state have close four momenta and are close to each other, that is, they have approximately the same coordinates.
As noted in Sec. 1, if dS or AdS symmetry works with good accuracy, then the operators $P_{\mu}$ are obtained from $M_{4\mu}$ by contraction. Therefore, it is natural to expect that for supersingletons in the bound state, the operators $M_{4\mu}$ have close eigenvalues. In particular, if $\Psi_{2}=\theta\Psi_{1}$ and for some value of $\mu$, $M_{4\mu}^{(1)}\Psi_{1}=\lambda_{1}\Psi_{1}$ then, as follows from Eqs. (8), $M_{4\mu}^{(2)}\Psi_{2}=\lambda_{1}\Psi_{2}$ and therefore, as follows from Eq. (10)
$M_{4\mu}(\Psi_{1}\times\Psi_{2})=2\lambda_{1}(\Psi_{1}\times\Psi_{2})$ (11)
Also, $\Psi_{1}$ and $\Psi_{2}=\theta\Psi_{1}$ can be treated as the states having the same coordinates. As discussed e.g., in [14], the position operators can be constructed from the operators $M_{\mu\nu}$. Therefore if for some values of $\mu$ and $\nu$, $M_{\mu\nu}^{(1)}\Psi_{1}=\lambda_{2}\Psi_{1}$ then, as follows from Eqs. (8), $M_{\mu\nu}^{(2)}\Psi_{2}=-\lambda_{2}\Psi_{2}$ and therefore, as follows from Eq. (10)
$M_{\mu\nu}(\Psi_{1}\times\Psi_{2})=0$ (12)
Now we have a natural explanation of the fact that, for a system consisting from supersingleton and antisupersingleton in a bound state, the eigenvalues of the operators $M_{4\mu}$ are much greater than the eigenvalues of the operators $M_{\mu\nu}$: as follows from Eqs. (8,11,12), the eigenvalues of operators $M_{4\mu}$ for individual supersingletons are included in the two-particle operators $M_{4\mu}$ with the same signs, while the eigenvalues of the operators $M_{\mu\nu}$ with different ones. Therefore, we have a natural explanation of the fact that for a particle consisting from supersingleton and antisupersingleton, Poincare symmetry works with high accuracy.
@@
Let us now consider the case when supersingletons entering a bound state are not antiparticles for each other. Then the IR of the AdS algebra for each supersingleton in a bound state is described by operators satisfying the relations (2) and there is no need to use the operator $\theta$. Even when we work in FQT and consider states of supersingletons in which the quantum numbers $j,k$ are much less than $p$, then, with high accuracy, we can apply standard mathematics. We assume that, although the numbers $j,k$ can be very large, they are still much less than $p$. Therefore, in what follows we consider Dirac supersingletons only within the framework of standard mathematics.
@@
We now treat $(d_{1}^{\prime},d_{2}^{\prime},d_{1}^{\prime\prime},d_{2}^{\prime\prime})$ as the operators in the Hilbert space related by Hermitian conjugation as $(d_{1}^{\prime})^{*}=d_{1}^{\prime\prime}$ and $(d_{2}^{\prime})^{*}=d_{2}^{\prime\prime}$. Then, as follows from Eq. (9), the norm squared of $e(j,k)$ equals
@@
and the normalized basis vectors can be defined asn
@@
In particle scattering experiments, the four-momenta and coordinates of initial and final particles are known with great accuracy. Therefore, these particles are described with good accuracy in the semiclassical approximation. In this case, the supersingleton wave functions $\sum_{jk}c(j,k){\tilde{e}}(j,k)$ are such that the coefficients $c(j,k)$ are not equal to zero only at $j\in(j_{1},j_{2}),\,\,k\in(k_{1},k_{2})$ where $j_{2}-j_{1}\ll j_{1},\,\,k_{2}-k_{1}\ll k_{1}$ and the values of $|c(j,k)|$ at such $j,k$ are approximately the same. We define the angular dependence of the coefficients as $c(j,k)=|c(j,k)|exp[i(j+k)\chi+i(j-k)\varphi]$. Then taking into account Eqs. (4,9,14) and the definition of the basis elements and the coefficients $c(j,k)$, direct calculation shows that, in semiclassical approximation, the operators $M_{ab}$ can be replaced by their numerical values:
@@
$M_{10}=jsin(2\varphi+2\chi)+ksin(2\varphi-2\chi),\,\,M_{04}=j+k+1$
@@
$M_{24}=jsin(2\varphi+2\chi)-ksin(2\varphi-2\chi)$ (15)
@@
As noted above, for a system of two free supersingletons 1 and 2, the AdS superalgebra representation for such a system is the tensor product of the representations for supersingletons 1 and 2, and the representation operators are the sums of the corresponding operators: $M_{ab}=M_{ab}^{(1)}+M_{ab}^{(2)}$.